# Turning Tabular Foundation Models into Graph Foundation Models

### v1

@@ 7 Conclusion
In this work, we have shown that tabular foundation models can be successfully employed for solving graph problems since they are able to process heterogeneous feature and target spaces. To show this, we proposed a simple G2T-FM framework, which converts a graph task into a tabular task by augmenting the initial features with graph information, and then applies a tabular foundation model. Our empirical results show the strong performance of G2T-FM, especially in comparison to other GFMs that fail to compete with well-tuned classic GNNs. This demonstrates that efficiently adapting to novel feature and target spaces is a very hard yet very important problem. Our approach is simple, but it provides significant advantages over prior attempts at developing GFMs and shows the potential of creating truly generalizable GFMs that can achieve strong results across diverse tasks and real-world applications of graph machine learning, regardless of the inherent feature and target spaces.
@@
Our proposed model, G2T-FM, currently serves as a proof of concept and has several noteworthy limitations that suggest future research directions.
First, the present version of G2T-FM uses only basic methods for processing graph structures. We plan to bring more graph-specific components into the framework, such as more complex aggregation mechanisms (including learnable and multi-hop aggregations) and cross-graph pretraining. We believe these extensions could enable the model to better capture graph-specific information and transfer knowledge across different graphs.
Second, because G2T-FM is based on TabPFNv2, it inherits several restrictions from this model. In particular, G2T-FM cannot handle classification datasets with more than 10 classes, and its training set size is limited to at most 10,000 samples. These limitations currently make it difficult to apply G2T-FM to large-scale datasets. As part of future work, we plan to experiment with other tabular foundation models within the G2T-FM framework to explore their potential benefits and limitations.
Appendix A Implementation Details
@@ A.1 G2T-FM
For the finetuning experiments, we follow the procedure outlined by Rubachev et al. (2025). Rather than using parameter-efficient finetuning, we opt for full model finetuning, as previous work indicates this yields better performance. We search for the optimal learning rate over a logarithmic grid of 10 values, ranging from $5\times 10^{-6}$ to $5\times 10^{-4}$.
@@
For the in-context learning experiments, we utilize a randomly initialized PEARL model whose weights are shared across all datasets. Interestingly, even without explicit training, this untrained PEARL model still produces useful representations for some datasets, as demonstrated in our ablation studies (Section 6). For the finetuning experiments, we jointly finetune PEARL and the TabPFNv2 backbone.
@@ A.2 GNNs and LightGBM
Our GNN setup closely follows the architecture and hyperparameter optimization procedure from GraphLand (Bazhenov et al., 2025), with two main differences. First, we introduce early stopping with a patience of 100 steps to accelerate training. Second, we use unified hyperparameter search spaces for both feature and target preprocessing, rather than dataset-specific spaces used in GraphLand. See our code for more details. Note that the last modification only affects the preprocessing hyperparameters, while the search grids for learning rate and dropout remain identical to those in GraphLand.
@@
Appendix B Equivariance
For graph problems, it is typically assumed that node IDs can be relabeled, feature columns can be reordered, and class IDs can be renamed without changing the underlying task. Hence, a model should not depend on these arbitrary choices. This motivates three symmetries for graph foundation models, as advocated in Finkelshtein et al. (2025): (i) feature permutation invariance; (ii) label permutation equivariance, and (iii) node permutation equivariance.
Formally, let $G$ be a group acting on inputs $\mathcal{X}$ and outputs $\mathcal{Y}$. A mapping $f:\mathcal{X}\to\mathcal{Y}$ is $G$-equivariant if $f(g\!\cdot\!x)=g\!\cdot\!f(x)$ for all $g\in G$, $x\in\mathcal{X}$. It is $G$-invariant if $f(g\!\cdot\!x)=f(x)$ for all $g$. In our context, relevant groups include node permutations $S_{|V|}$, feature (column) permutations $S_{d}$, and label permutations $S_{|C|}$. Here, $S_{n}$ denotes the symmetric group on a set of $n$ elements (i.e., the group of all permutations), $|V|$ is the number of nodes, $d$ is the number of features, and $|C|$ stands for the number of classes.
@@ Appendix B Equivariance
Our model, G2T-FM, uses TabPFNv2 as its backbone. It is therefore important to analyze its symmetry properties. TabPFNv2 is sample permutation equivariant and feature permutation invariant in distribution. This property holds because the model applies random positional encodings to its input features. As these encodings are sampled independently and identically, the output distribution is unaffected by the order of the feature columns. By default, TabPFNv2 is not label permutation equivariant. However, this can be achieved with a simple modification that we incorporate into our implementation. Specifically, during each forward pass, we randomly permute the ordinal encodings assigned to the class labels. This procedure ensures that the model becomes label permutation equivariant in distribution.
G2T-FM
Besides TabPFNv2, G2T-FM also employs several other components, including NFA, PEARL, and simple structure-based features (degree, pagerank, Laplacian eigenvectors). NFA is node- and feature-equivariant. Structural features such as node degree, PageRank, and Laplacian eigenvectors are all node-equivariant by construction. The PEARL framework, which we also use, is node-equivariant in distribution. The combination of these components makes G2T-FM node permutation equivariant in distribution. Overall, in distribution, G2T-FM has all the desired symmetries.

### v2

@@ 7 Conclusion
In this work, we have shown that tabular foundation models can be successfully employed for solving graph problems since they are able to process heterogeneous feature and target spaces. To show this, we proposed a simple G2T-FM framework, which converts a graph task into a tabular task by augmenting the initial features with graph information, and then applies a tabular foundation model. Our empirical results show the strong performance of G2T-FM, both in the in-context and finetuning regimes. In particular, G2T-LimiX outperforms all well-tuned GNN baselines in terms of the average rank even in the in-context learning regime. After finetuning, G2T-LimiX achieves the best results on most of the datasets, often outperforming the best GNN by a significant margin. Our approach is simple, but it provides significant advantages over prior attempts at developing GFMs and shows the potential of creating truly generalizable GFMs that can achieve strong results across diverse tasks and real-world applications of graph machine learning, regardless of the inherent feature and target spaces. Importantly, G2T-FM can be combined with any tabular foundation model. Thus, future advancements in tabular foundation models can be easily transferred to the graph ML domain.
@@
Despite its strong performance, the proposed G2T-FM framework is only a first step towards utilizing models and ideas from the tabular domain for developing truly generalizable GFMs. Hence, our work has several noteworthy limitations that suggest future research directions.
First, the present version of G2T-FM uses only basic methods for processing graph structures. Future research can bring more graph-specific components into the framework, such as more complex aggregation mechanisms (including learnable and multi-hop aggregations) and cross-graph pretraining. We believe these extensions could enable the model to better capture graph-specific information and transfer knowledge across different graphs.
Second, G2T-FM inherits several restrictions from its TFM backbone. In particular, when endowed with TabPFNv2, G2T cannot handle classification datasets with more than 10 classes, and its training set size is limited to at most 10,000 samples. These limitations currently make it difficult to apply G2T-FM to large-scale datasets. However, future research on tabular foundation models can alleviate these limitations and the new TFMs can be directly used within the G2T-FM framework.
Appendix A Concurrent work
Soon after we released the first version of this preprint on arXiv, an independent concurrent work (Hayler et al., 2025) also explored using tabular foundation models for graph tasks. Our work differs in several important ways. First, we focus on real-world datasets from the recently introduced GraphLand benchmark (Bazhenov et al., 2025) which contains a variety of classification and regression tasks covering more realistic problems compared to classic graph benchmarks (Bechler-Speicher et al., 2025). Second, we evaluate TFMs in both in-context learning and finetuning regimes, and we show that finetuning substantially improves performance and, on average, enables TFMs to outperform strong baselines. In this revision of the paper, we have also added a more recent tabular foundation model, LimiX, allowing us to improve the performance of G2T-FM even further.
Appendix B Implementation Details
@@ B.1 G2T-FM
For the finetuning experiments, we follow the procedure outlined by Rubachev et al. (2025). Rather than using parameter-efficient finetuning, we opt for full model finetuning, as previous work indicates this yields better performance. We search for the optimal learning rate over the logarithmic grid of 10 values, ranging from $5\times 10^{-6}$ to $5\times 10^{-4}$.
@@
For the in-context learning experiments, we utilize a randomly initialized PEARL model whose weights are shared across all datasets. Interestingly, even without explicit training, this untrained PEARL model still produces useful representations for some datasets, as demonstrated in our ablation studies (Section 6). For the finetuning experiments, we jointly finetune PEARL and the TFM backbone.
@@ B.2 GNNs and LightGBM
Our GNN setup closely follows the architecture and hyperparameter optimization procedure from GraphLand (Bazhenov et al., 2025), with two main differences. First, we introduce early stopping with a patience of 100 steps to accelerate training. Second, we use unified hyperparameter search spaces for both feature and target preprocessing, rather than dataset-specific spaces used in GraphLand. See our code for more details. Note that this latter modification only affects the preprocessing hyperparameters, while the search grids for learning rate and dropout remain identical to those in GraphLand.
@@
Appendix C Symmetries: equivariance and invariance
For graph problems, it is typically assumed that node IDs can be relabeled, feature columns can be reordered, and class IDs can be renamed without changing the underlying task. Hence, a model should not depend on these arbitrary choices. This motivates three symmetries for graph foundation models, as advocated in Finkelshtein et al. (2025): (i) feature permutation invariance; (ii) label permutation equivariance; and (iii) node permutation equivariance.
Formally, let $G$ be a group acting on inputs $\mathcal{X}$ and outputs $\mathcal{Y}$. A mapping $f:\mathcal{X}\to\mathcal{Y}$ is $G$-equivariant if $f(g\!\cdot\!x)=g\!\cdot\!f(x)$ for all $g\in G$, $x\in\mathcal{X}$. It is $G$-invariant if $f(g\!\cdot\!x)=f(x)$ for all $g$. In our context, relevant groups include node permutations $S_{|V|}$, feature (column) permutations $S_{d}$, and label permutations $S_{|C|}$. Here, $S_{n}$ denotes the symmetric group on a set of $n$ elements (i.e., the group of all permutations), $|V|$ is the number of nodes, $d$ is the number of features, and $|C|$ is the number of classes.
@@ Appendix C Symmetries: equivariance and invariance
Below, we discuss the symmetries of G2T-FM. In particular, we show that all components added to TFM are equivariant in distribution, so the resulting symmetries of G2T-FM depend on the symmetries of the TFM backbone. As an example, we analyze the symmetries of TabPFNv2.
G2T-FM components
The symmetries of G2T-FM rely on the symmetries of the chosen TFM. Let us show that if the TFM has feature permutation invariance, label permutation equivariance and sample permutation equivariance (in distribution), then G2T-FM also has all the desired symmetries. Indeed, G2T-FM employs several components: NFA, PEARL, and simple structure-based features (degree, PageRank, Laplacian eigenvectors). NFA is node- and feature-equivariant. Structural features such as node degree, PageRank, and Laplacian eigenvectors are all node-equivariant by construction. The PEARL framework, which we also use, is node-equivariant in distribution. The combination of these components makes G2T-FM node permutation equivariant in distribution. Thus, all the desired symmetries are preserved.
@@
TabPFNv2 is sample permutation equivariant and feature permutation invariant in distribution. This property holds because the model applies random positional encodings to its input features. As these encodings are sampled independently and identically, the output distribution is unaffected by the order of the feature columns. By default, TabPFNv2 is not label permutation equivariant. However, this can be achieved with a simple modification that we incorporate into our implementation. Specifically, during each forward pass, we randomly permute the ordinal encodings assigned to the class labels. This procedure ensures that the model becomes label permutation equivariant in distribution.